OpenStudy (anonymous):

Let F(x)= f(f(x)) and G(x)=(F(x))^2. You also know that f(4)=3, f(3)=3, f'(3)=14, f'(4)=3. Find F'(4) and G'(4).

OpenStudy (anonymous):

Use the chain rule: \[ F'(x) = f'(f(x))\cdot f'(x) \\ G'(x) = 2F(x) \cdot F'(x) = 2f(f(x)) \cdot f'(f(x)) \cdot f'(x) \]

OpenStudy (anonymous):

@youngbuck2123 got the idea?

OpenStudy (anonymous):

I think so, when computed do you come up with 252 as the answer??

OpenStudy (anonymous):

I don't know the answer cause I haven't done it. If you show your work I can check it though.

OpenStudy (anonymous):

answer came out correct, thanks for your help